OpenStudy (anonymous):

Solve for x: −5|x + 1| = 10 x = 0 x = −3 and x = 1 x = −1 and x = 3 No solution

OpenStudy (nurali):

−5|x + 1| = 10 Divide by -5 |x + 1|=-2 x+1=-2 or x+1=-(-2) x+1=-2 or x+1=2 x=-2-1 or x=2-1 x=-3 or x=1

OpenStudy (anonymous):

whats the answer?

OpenStudy (nurali):

Second option